A rectangular sheet of wood has four  small squares removed from the corners. It is then cut to make a box that is 5 cm by 4 cm with a volume of 60 cm 3 . (Four pieces of size A4 are removed.) Find the original area of the sheet of wood.

Grade:10

1 Answers

Aman Bansal
12 years ago

Dear Sathyaram,

 l x h x w = 60 
     h(5)(4) = 60
     h = 3

     A(1) = 4 x 5 = 20
     2A(2) = 2(3)(5) = 30
     2A(3) = 2(3)(4) = 24
     4A(4) = 4(3)(3) = 36

     Total A = A(1) + 2A(2) + 2A(3) + 4A(4)
             = 20 + 30 + 24 + 36
             = 110cm2

Best Of Luck...!!!!

Cracking IIT just got more exciting,It’s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and winexciting gifts by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple click here to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert